1.  A boy weighing 150 N goes up and down a flight of stairs. The vertical distance between the floor is 5.5 m. How much work is done by gravity?



1
Expert's answer
2021-11-19T10:51:00-0500

(a) Let's find the work done by the gravity in case the boy goes up a flight of stairs:


"W=Fdcos\theta=W_{boy}dcos\theta,""W=150\ N\times5.5\ m\times cos180^{\circ}=-825\ J."


The sign minus means that the work done against the direction of gravity.

(b) Let's find the work done by the gravity in case the boy goes down a flight of stairs:


"W=Fdcos\theta=W_{boy}dcos\theta,""W=150\ N\times5.5\ m\times cos0^{\circ}=825\ J."


The sign plus means that the work done in the direction of gravity.